# English Please: Evaluating Machine Translation for Multilingual Bug Reports

### v2

@@ Abstract
Accurate translation of bug reports is critical for efficient collaboration in global software development. In this study, we conduct the first comprehensive evaluation of machine translation (MT) performance on bug reports, analyzing the capabilities of DeepL, AWS Translate, and ChatGPT using data from the Visual Studio Code GitHub repository, specifically focusing on reports labeled with the english-please tag. To thoroughly assess the accuracy and effectiveness of each system, we employ multiple machine translation metrics, including BLEU, BERTScore, COMET, METEOR, and ROUGE. Our findings indicate that DeepL consistently outperforms the other systems across most automatic metrics, demonstrating strong lexical and semantic alignment. AWS Translate performs competitively, particularly in METEOR, while ChatGPT lags in key metrics. This study underscores the importance of domain adaptation for translating technical texts and offers guidance for integrating automated translation into bug-triaging workflows. Moreover, our results establish a foundation for future research to refine machine translation solutions for specialized engineering contexts. The code and dataset for this paper are available at GitHub -https://github.com/av9ash/English-Please.
@@
Multilingual Bug Reports, Machine Translation (MT), ChatGPT, AWS Translate, DeepL, Visual Studio Code, english-please, BLEU, BERT Score, COMET, METEOR, ROUGE, Software Development, Domain-Specific Translation, Technical Text Localization, Large Language Models
@@ I Introduction
As global software development increasingly depends on multilingual collaboration, the accurate translation of technical documents has become a crucial challenge. Among these, bug reports present unique difficulties due to using domain-specific jargon, software version references, and embedded code snippets—elements that often challenge generic machine translation (MT) systems [1]. While recent advances in neural MT have significantly improved translation fluency and accuracy [2, 3], their effectiveness on specialized technical texts, such as software bug reports, remains an open question.
@@
We evaluate the translation performance of three leading MT systems: ChatGPT, AWS Translate, and DeepL. Each system translates non-English bug reports into English. We assess their outputs using five established quality metrics: BLEU [6], BERTScore [7], COMET [8], METEOR [9], and ROUGE [10].
By leveraging these diverse metrics, we aim to understand translation quality comprehensively. Specifically, this paper addresses the following research questions:
@@
  • RQ2: How do ChatGPT, AWS Translate, and DeepL compare in preserving the technical precision of VS Code bug reports?
By answering these questions, we provide insights into the strengths and limitations of modern MT systems in software development contexts. Our findings will help researchers and practitioners optimize the integration of automated translation into bug-triaging workflows.
The remainder of this paper is structured as follows: Section II contextualizes our work within the existing literature, Section III describes our dataset and preprocessing approach, Section IV details our experimental methodology, Section V presents results and analysis, Section VI discusses implications and limitations, and Section VII concludes with final remarks and future directions.
@@ II Related Work
Finally, while industry-driven solutions such as AWS Translate, DeepL, and ChatGPT are commonly employed for multilingual documentation and user support, published comparative analyses specific to software bug reports remain limited. Existing studies typically focus on localization scenarios or general test suites that may not capture the nuanced requirements of bug triaging and software maintenance. As such, a deeper examination of how these systems handle authentic bug reports is necessary and timely. Our work aims to fill this gap by systematically evaluating these three MT solutions on bug reports from the Visual Studio Code repository, offering insights into their strengths and limitations in real-world software engineering.
@@ III Dataset
The dataset used in this study comprises bug reports extracted from the Visual Studio Code GitHub repository [4], specifically focusing on those labeled with the english-please tag. This label indicates that the original submission was written in a language other than English, providing a clear signal for multilingual content. We collected issues spanning five years (March 2019–June 2024), as represented in Figure 1, to ensure a sufficient variety of bug types, user environments, and technical contexts.
To facilitate a reliable evaluation, we applied several filtering steps:
  1. 1.
    Issue Validity: We excluded duplicates and issue threads lacking a clear bug description or reproduction steps. Furthermore, Figure 2 illustrates the top 20 labels in the dataset.
  2. 2.
    Language Scope: Our focus was on non-English reports to align with the english-please label, ensuring the dataset authentically represented multilingual content. Additionally, Table I provides a detailed breakdown of the language distribution of bug reports, highlighting the linguistic diversity within our dataset.
  3. 3.
    Content Cleaning: We removed sensitive user data, standardized code snippets, and retained only the textual elements necessary for translating bug descriptions, error messages, and steps to reproduce.
Each curated issue was paired with a reference English translation verified by one or more bilingual contributors. This yielded a final corpus of 1,300 multilingual bug reports, reflecting a broad spectrum of functional, UI, and performance-related issues. Bug reports can vary substantially in their level of technical detail and specialized terminology. Our goal was to capture this diversity to provide a robust test set for evaluating the translation performance of ChatGPT, AWS Translate, and DeepL.
@@ IV Methodology
This section describes the experimental design used to compare the performance of ChatGPT, AWS Translate, and DeepL on the multilingual bug reports presented in Section 2. Our methodology encompasses the data collection and preparation pipeline, the translation process for each system, and the evaluation metrics employed to gauge translation quality.
IV-A Experimental Setup
We followed a systematic process aligned with prior MT evaluation studies [11, 12] to ensure reproducibility and reliability:
@@ IV-A Experimental Setup
    Preprocessing: Consistent formatting is crucial when comparing outputs from different MT engines [5]. We standardized the bug report structure by removing extraneous metadata (e.g., GitHub labels and user mentions) and retaining only text relevant to problem descriptions, error messages, and reproduction steps.
@@
    Annotation Protocol: Each bug report in the test set was accompanied by a reference English translation validated by bilingual reviewers. These reference translations serve as ground truth for our evaluation metrics [12].
IV-B Translation Systems
We evaluated three MT services:
  1. 1.
    ChatGPT: We accessed ChatGPT through its API, ensuring uniform prompts to elicit translations. Recent neural architectures [3] underpin large language models like ChatGPT, enabling it to handle various linguistic phenomena. However, ChatGPT’s internal model updates can lead to temporal variations in performance.
  2. 2.
    AWS Translate: This commercial solution offers a ready-to-use platform for general-domain translations. While AWS Translate has demonstrated strong performance on conventional text, it may face domain adaptation challenges for bug-report content [11].
@@ IV-B Translation Systems
    DeepL: DeepL’s core engine leverages Transformer-based methods and is reputed for its robust handling of European languages. Nevertheless, the translation of domain-specific terms (e.g., method names, code snippets) remains uncertain in specialized contexts [5].
@@ V-A Automatic Metrics
Table II summarizes the results of the five automatic evaluation metrics—BLEU, METEOR, ROUGE, BERTScore, and COMET—calculated over 1,300 multilingual bug reports. We highlighted the highest scores for each metric in bold.
@@
Metric (Mean) ChatGPT DeepL AWS Translate
BLEU 0.42 0.53 0.49
METEOR 0.60 0.59 0.62
COMET 0.23 0.32 0.32
BERTScore (F1) 0.73 0.76 0.75
ROUGE-L (F1) 0.64 0.68 0.67
BLEU: DeepL achieved the highest BLEU score (0.53), followed by AWS Translate (0.49), with ChatGPT scoring the lowest (0.42), suggesting that DeepL had the most substantial n-gram overlap with the reference translations. However, BLEU alone may not fully capture nuanced meaning in technical text.
METEOR: AWS Translate led with a score of 0.62, outperforming ChatGPT (0.60) and DeepL (0.59), indicating that AWS Translate better aligns with reference translations in terms of synonyms, stems, and paraphrases.
ROUGE: DeepL achieved the highest ROUGE-L F1 score (0.68), followed closely by AWS Translate (0.67) and ChatGPT (0.64). Since bug reports often contain specific error messages and repeated domain terms, a higher ROUGE score may indicate a tendency toward literal phrasing.
BERTScore: DeepL attained the highest BERTScore F1 (0.76), suggesting strong semantic alignment with reference translations. AWS Translate followed with 0.75, while ChatGPT scored the lowest at 0.73, highlighting minor differences in capturing contextual meaning across bug descriptions.
COMET: DeepL and AWS Translate achieved the highest COMET score (0.32), with ChatGPT trailing at 0.23. Since COMET relies on pre-trained language models to assess semantic fidelity, these results suggest that DeepL and AWS Translate maintained firmer semantic consistency in domain-specific text such as bug reports.
Accuracy: Table III compares the three translation models—ChatGPT, DeepL, and AWS Translate—for source language identification. While AWS Translate achieved the highest accuracy (0.4786), ChatGPT performed better in precision (0.7518), suggesting a higher proportion of correct language identifications among the predicted instances. DeepL, while achieving moderate performance across all metrics, showed a balance between recall (0.6958) and F1-score (0.6634). The confusion matrix is visualized in Figure 4(c).
@@
Metric ChatGPT DeepL AWS Translate
Accuracy 0.415 0.464 0.479
Precision 0.752 0.651 0.712
Recall 0.760 0.696 0.786
F1-score 0.717 0.663 0.737
Figure 4: Confusion matrices for language identification by AWS Translate, DeepL, and ChatGPT.
@@
(b) DeepL
(c) ChatGPT
V-B Summary of Findings
DeepL outperformed ChatGPT and AWS Translate across most translation metrics, particularly BLEU, ROUGE, BERTScore, and COMET, demonstrating superior accuracy and semantic alignment. AWS Translate achieved the highest METEOR score, indicating more vigorous handling of synonyms, stemming, and paraphrasing. ChatGPT exhibited the weakest performance across all metrics, reflecting lower n-gram overlap and reduced semantic consistency.
No single model consistently outperformed the others across all evaluation criteria for language identification. The optimal choice depends on whether precision, recall or overall accuracy is prioritized for multilingual translation tasks.
VI Discussion
The results presented in Section V illustrate both the promise and limitations of applying leading MT systems to the specialized domain of bug reports. DeepL’s strong performance across most automatic metrics underscores its effectiveness in handling structured text and maintaining semantic consistency. AWS Translate also demonstrated strengths, particularly in METEOR, while ChatGPT lagged in multiple metrics. However, as noted by Chu and Wang [1], domain adaptation remains a significant challenge for neural MT systems, especially when handling specialized tokens such as code references or UI labels. These findings highlight the need for further MT model refinement to accommodate domain-specific terminology in software engineering tasks better.
VI-A Interpreting Translation Metrics
The discrepancies among BLEU, METEOR, and ROUGE scores highlight the inherent limitations of each metric in capturing the full spectrum of translation quality [12]. For instance, DeepL’s higher BLEU and ROUGE scores indicate strong lexical and structural alignment, yet ChatGPT’s lower BERTScore and COMET results suggest weaker semantic consistency. Interestingly, AWS Translate led in METEOR but did not achieve the highest scores in other semantic metrics, suggesting strengths in lexical variation but challenges in deeper contextual alignment. These findings align with prior observations that purely $n$-gram-based metrics can fail to capture contextual nuances in technical text [5]. By contrast, metrics like BERTScore and COMET, where DeepL and AWS Translate outperformed ChatGPT, leverage contextual embeddings to provide a richer view of semantic fidelity, an approach increasingly recommended for domain-specific evaluations [14, 15].
VI-B Contextual Limitations and Model Updates
One noteworthy caveat involves the evolving nature of cloud-based MT systems. ChatGPT, for instance, is updated regularly, potentially causing temporal variations in translation performance. AWS Translate and DeepL also deploy algorithmic improvements over time. As with prior studies on continuously evolving MT frameworks [12], our findings represent a snapshot of performance rather than a static benchmark. Future research should consider conducting repeated evaluations over extended periods to capture longitudinal trends better.
VI-C Implications for Software Engineering
Our study highlights the necessity of domain-specific adaptation when translating bug reports, echoing the broader literature on specialized MT tasks [11]. Software teams may benefit from integrating glossaries or custom dictionaries into translation pipelines. Such resources could address recurring terminology (e.g., function names, proprietary module references) that appear throughout a codebase. Tools that automatically identify and preserve code tokens could also mitigate the risk of incorrect “translations” of critical technical elements.
VI-D Limitations
Although our dataset includes a representative variety of bug reports from the Visual Studio Code repository, it focuses on a single project with its ecosystem of terminology and user community. Expanding the dataset to include bug reports from multiple software systems with diverse technology stacks could assess wider applicability. Moreover, we primarily relied on bilingual software engineers for initial labeling; while their expertise adds domain knowledge, cross-lingual proficiency levels, and personal biases can still influence subjective judgments.
VI-E Summary
Our findings emphasize that while the latest MT services can produce increasingly fluent and contextually coherent translations, they still exhibit non-trivial deficiencies in domain adequacy. Bridging this gap requires a combination of advanced neural architectures, specialized training approaches, and the integration of software-specific resources. This work thus contributes to a clearer understanding of the interplay between general-purpose MT systems and the demands of real-world bug reporting scenarios, setting the stage for future improvements and cross-disciplinary collaborations.
VII Conclusion and Future Work
This paper has examined the capability of three leading machine translation (MT) services—DeepL, AWS Translate, and ChatGPT—to accurately translate multilingual bug reports sourced from the Visual Studio Code repository. Our findings indicate that AWS Translate performs best in language identification (RQ1), achieving the highest accuracy (0.479) and F1-score (0.737), followed by DeepL and ChatGPT, suggesting that AWS Translate is more effective in detecting source languages, which is crucial for efficient translation workflows.
In preserving the technical precision of VS Code bug reports (RQ2), DeepL consistently outperforms the other systems across most automatic metrics, including BLEU, ROUGE, and BERTScore, demonstrating strong lexical and semantic fidelity. AWS Translate performs competitively, particularly in METEOR, suggesting it handles paraphrases well, while ChatGPT lags in most metrics, indicating weaker alignment with the original bug report content.
VII-A Key Observations
  • Language Identification Performance (RQ1): AWS Translate achieved the highest accuracy and F1-score, suggesting it is the most reliable source language detecting. DeepL followed closely, while ChatGPT had the lowest identification accuracy.
  • Lexical vs. Semantic Accuracy (RQ2): DeepL achieved the highest BLEU, ROUGE, and BERTScore results, indicating strong preservation of technical precision in bug reports. AWS Translate performed best in METEOR, highlighting its ability to handle paraphrases, while ChatGPT lagged in most metrics.
  • Metric Variability: The discrepancies across BLEU, METEOR, ROUGE, BERTScore, and COMET emphasize the importance of multi-metric evaluations. While n-gram-based metrics capture lexical matches, embedding-based approaches such as BERTScore and COMET provide a deeper assessment of semantic alignment [14, 15].
VII-B Future Directions
Several avenues warrant further investigation:
  1. 1.
    Custom Domain Adapters: Building on existing work [11], future research could explore specialized glossaries or fine-tuned models to mitigate the loss of critical technical details, involving integrating bug report ontologies or code-aware tokenization methods into the translation pipeline.
  2. 2.
    Longitudinal Analysis: Cloud-based MT services are frequently updated, potentially altering translation quality over time [12]. Periodic re-evaluations could reveal performance trends and help practitioners decide when to re-translate or re-verify critical documentation.
  3. 3.
    Broader Dataset Coverage: Although this study focused on VS Code bug reports, expanding to other large-scale projects across different programming languages and technology stacks would yield a more generalizable understanding of MT performance in software engineering contexts.
In conclusion, while modern MT services have made significant strides, particularly in fluency and general-domain accuracy, further adaptation is required to meet the nuanced needs of multilingual bug reporting fully. This paper provides a framework for evaluating domain-specific translation tasks. We hope these findings will spur further innovations in integrating MT within software engineering pipelines, ultimately enabling more seamless and efficient global collaboration.

### v3

# English Please: Evaluating Machine Translation with Large Language Models for Multilingual Bug Reports

@@ Abstract
Accurate translation of bug reports is critical for efficient collaboration in global software development. In this study, we conduct the first comprehensive evaluation of machine translation (MT) performance on bug reports, analyzing the capabilities of DeepL, AWS Translate, and large language models such as ChatGPT, Claude, Gemini, LLaMA, and Mistral using data from the Visual Studio Code GitHub repository, specifically focusing on reports labeled with the english-please tag. To assess both translation quality and source language identification accuracy, we employ a range of MT evaluation metrics—including BLEU, BERTScore, COMET, METEOR, and ROUGE—alongside classification metrics such as accuracy, precision, recall, and F1-score. Our findings reveal that while ChatGPT (gpt-4o) excels in semantic and lexical translation quality, it does not lead in source language identification. Claude and Mistral achieve the highest F1-scores (0.7182 and 0.7142, respectively), and Gemini records the best precision (0.7414). AWS Translate shows the highest accuracy (0.4717) in identifying source languages. These results highlight that no single system dominates across all tasks, reinforcing the importance of task-specific evaluations. This study underscores the need for domain adaptation when translating technical content and provides actionable insights for integrating MT into bug-triaging workflows. The code and dataset for this paper are available at GitHub—https://github.com/av9ash/English-Please.
@@
Multilingual Bug Reports, Machine Translation (MT), ChatGPT, Claude, Gemini, LLaMA, Mistral, AWS Translate, DeepL, Visual Studio Code, english-please, BLEU, BERTScore, COMET, METEOR, ROUGE, Software Development, Domain-Specific Translation, Technical Text Localization, Large Language Models, Source Language Identification
@@ I Introduction
As global software development increasingly depends on multilingual collaboration, the accurate translation of technical documents has become a crucial challenge. Among these, bug reports present unique difficulties due to domain-specific jargon, software version references, and embedded code snippets—elements that often challenge generic machine translation (MT) systems [1]. While recent advances in neural MT have significantly improved translation fluency and accuracy [2, 3], their effectiveness on specialized technical texts, such as software bug reports, remains an open question.
@@
We evaluate the translation and language identification performance of several leading MT systems: ChatGPT, AWS Translate, DeepL, Claude, Gemini, Mistral, and LLaMA. Each system is tasked with identifying the source language and translating non-English bug reports into English. Language identification accuracy is especially critical, as an incorrect assumption about the source language can cascade into flawed translations. Our evaluation of language detection is based on four established metrics: accuracy, precision, recall, and F1-score.
To evaluate translation quality, we use five complementary metrics: BLEU[6], BERTScore[7], COMET[8], METEOR[9], and ROUGE [10]. By leveraging these diverse measures, we aim to assess not only fluency and adequacy but also the preservation of technical content.
Specifically, this paper addresses the following research questions:
@@
  • RQ2: How do state-of-the-art machine translation (MT) models—such as Amazon Translate and DeepL—compare with large language models (LLMs) like ChatGPT, Claude, Gemini, LLaMA, and Mistral in preserving the technical precision of Visual Studio Code (VS Code) bug reports?
By answering these questions, we provide insights into the strengths and limitations of modern MT systems in software development contexts. Our findings aim to inform researchers and practitioners seeking to integrate automated translation into multilingual bug-triaging workflows effectively.
The remainder of this paper is structured as follows: Section II contextualizes our work within the existing literature, Section III describes our dataset, Section IV details our experimental methodology, Section V presents results and analysis, and Section VI offers concluding remarks and discussion.
@@ II Related Work
Finally, while industry-driven solutions such as AWS Translate, DeepL, and LLMs are commonly employed for multilingual documentation and user support, published comparative analyses specific to software bug reports remain limited. As such, a deeper examination of how these systems handle authentic bug reports is necessary and timely. Our work aims to fill this gap by systematically evaluating these MT solutions on bug reports from the Visual Studio Code repository, offering insights into their strengths and limitations in real-world software engineering.
@@ III Dataset
The dataset used in this study comprises bug reports extracted from the Visual Studio Code GitHub repository [4], specifically those labeled english-please—a tag typically indicating submissions originally written in a language other than English, and thus serving as a reliable marker for multilingual content. Spanning five years (March 2019–June 2024), the dataset captures a wide variety of bug types, user environments, and technical contexts, as illustrated in Figure 1. Although most english-please issues were indeed non-English, some were already in English; these were retained to reflect real-world labeling inconsistencies and maintain dataset authenticity, as shown in Figure 2. The study focuses primarily on non-English reports, consistent with the tag’s intended purpose, to accurately represent the multilingual challenges in developer communication. The dataset’s linguistic diversity is summarized in Table I.
Each selected issue was paired with a reference English translation, verified by one or more bilingual contributors. This process yielded a final corpus of 1,300 multilingual bug reports covering a broad range of functional, UI, and performance-related issues. Given that bug reports vary widely in technical detail and domain-specific language, the dataset was curated to preserve this diversity and provide a robust testbed for evaluating the translation performance of all models considered.
@@ IV Methodology
This section outlines our approach to evaluating the translation and language identification performance of seven state-of-the-art machine translation (MT) models on a curated dataset of 1,300 multilingual bug reports. We tested both commercial services (AWS Translate, DeepL, ChatGPT) and research-grade LLMs (Claude, Gemini, Mistral, LLaMA) under consistent conditions. Our evaluation combines translation quality metrics—BLEU, METEOR, ROUGE, BERTScore, and COMET—with classification measures such as accuracy, precision, recall, F1-score, and confusion matrix analysis. This methodology enables a comprehensive comparison of how effectively each model handles both technical translation and source language detection in a multilingual software development setting.
IV-A Translation Systems
We evaluated seven state-of-the-art machine translation (MT) models, encompassing both commercial services and open-source large language models (LLMs):
@@ IV-A Translation Systems
    AWS Translate: A production-grade MT service, AWS Translate offers scalable, general-purpose translations. It performs reliably on standard text but may encounter challenges when translating domain-specific content such as developer jargon, inline code, or structured formatting [11].
@@
    DeepL: DeepL is widely recognized for producing fluent and coherent translations, particularly for European languages. It leverages a Transformer-based architecture optimized for natural language generation. However, it may inconsistently handle technical constructs such as method names, variable identifiers, and embedded code [5].
@@
    ChatGPT (GPT-4o): Accessed via the OpenAI API, ChatGPT is a general-purpose large language model built on Transformer-based architectures [3]. It supports multilingual translation through prompt-based interaction and is capable of handling complex linguistic phenomena. However, performance may vary due to periodic updates to the underlying model.
  4. 4.
    Claude (Sonnet 3.7): Developed by Anthropic, Claude is an LLM designed with a focus on safety and contextual understanding. It demonstrates strong multilingual capabilities and emphasizes the preservation of meaning, which is particularly beneficial for maintaining the intent of technical descriptions in bug reports.
  5. 5.
    Gemini (Pro 1.5): Gemini (formerly Bard), developed by Google DeepMind, combines multilingual translation abilities with large-scale knowledge integration. It excels in precision during language identification and is capable of responding to natural-language prompts with detailed translations.
  6. 6.
    Mistral (Pixtral Large): Mistral is an open-weight, decoder-only Transformer model engineered for high efficiency and multilingual performance. It demonstrates strong recall and F1-scores in language identification tasks, suggesting robust coverage across diverse linguistic patterns.
  7. 7.
    LLaMA (Maverick): LLaMA (Large Language Model Meta AI), developed by Meta, is a family of open-source foundation models trained on diverse multilingual corpora. Although not explicitly optimized for translation tasks, it performs competitively in multilingual settings due to its extensive pretraining on varied textual data.
@@ V-A Automatic Metrics
Table II summarizes the results of five automatic evaluation metrics—BLEU, METEOR, ROUGE, BERTScore, and COMET—computed over 1,300 multilingual bug reports. The highest scores for each metric are highlighted in bold, and models are listed in descending order of overall performance.
@@
Model BERTScore BLEU COMET METEOR ROUGE
GPT 0.7838 49.32 0.3636 0.6613 0.7128
Mistral 0.7597 45.85 0.3325 0.5895 0.6903
Claude 0.7626 43.68 0.3417 0.5807 0.6861
Gemini 0.7595 43.87 0.3604 0.5721 0.6854
DeepL 0.7589 43.06 0.2289 0.5570 0.6680
Llama 0.7576 39.42 0.2313 0.5923 0.6661
AWS 0.7487 41.90 0.2288 0.5792 0.6599
BLEU: ChatGPT (gpt-4o) achieved the highest BLEU score (49.32), followed by Mistral (45.85) and Gemini (43.87), indicating that ChatGPT had the greatest n-gram overlap with the reference translations. This suggests strong surface-level alignment, though BLEU may still fall short in capturing deeper semantic equivalence, particularly in technical text.
METEOR: ChatGPT also led in METEOR (0.6613), outperforming LLaMA (0.5923) and Mistral (0.5895), suggesting superior handling of synonyms, morphological variants, and paraphrases. This indicates that ChatGPT’s outputs were more aligned with the reference in terms of linguistic diversity and fluency.
ROUGE: ChatGPT recorded the highest ROUGE-L F1 score (0.7128), followed by Mistral (0.6903) and Claude (0.6861). Given the technical nature of bug reports, where exact phrasing often recurs, this result reflects ChatGPT’s strength in preserving key terminologies and phrasings.
BERTScore: ChatGPT obtained the highest BERTScore (0.7838), indicating the strongest semantic alignment with the reference translations. Claude (0.7626) and Gemini (0.7595) followed closely, further supporting ChatGPT’s contextual accuracy in domain-specific content.
COMET: ChatGPT again topped the scores with 0.3636, followed by Gemini (0.3604) and Claude (0.3417). As COMET leverages pretrained models for semantic assessment, these scores reinforce ChatGPT’s ability to preserve meaning in complex, technical inputs.
V-B Source Language Identification
Table III compares the language identification performance of seven translation models. AWS Translate achieved the highest accuracy (0.4717), indicating its strength in correctly identifying the source language. However, Gemini exhibited the highest precision (0.7414), meaning it was most likely to be correct when it did make a prediction. Claude led in recall (0.7581), capturing the largest proportion of actual positive instances, and also achieved the highest F1-score (0.7182), reflecting a strong balance between precision and recall. Among all models, GPT, Mistral, LLaMA, and Gemini demonstrated comparably high precision and recall, highlighting their consistency across predictions. DeepL maintained a balanced performance across all metrics, with an F1-score of 0.6716. The confusion matrix is visualized in Figure LABEL:fig:cm_plots.
@@ V-B Source Language Identification
Model Accuracy Precision Recall F1-score
AWS 0.4717 0.6788 0.7488 0.7027
DeepL 0.4586 0.6598 0.6954 0.6716
Claude 0.4079 0.7129 0.7581 0.7182
GPT 0.4070 0.7130 0.7190 0.6788
Mistral 0.4062 0.7120 0.7574 0.7142
LLaMA 0.4029 0.7098 0.7557 0.7120
Gemini 0.4005 0.7414 0.7639 0.7123
@@
(b) Claude
(c) DeepL
(d) Gemini
(e) ChatGPT
(f) Llama
(g) Mistral
Figure 4: Confusion matrices for language identification across seven machine translation models. Darker diagonal entries indicate correct classifications.
V-C Summary of Findings
ChatGPT (gpt-4o) outperformed all other models across all five automatic translation metrics—BLEU, METEOR, ROUGE, BERTScore, and COMET—demonstrating strong surface-level accuracy and deep semantic alignment in multilingual bug report translations. Mistral and Gemini also performed competitively, particularly in BLEU and COMET, while Claude showed strong performance in BERTScore and ROUGE. DeepL and AWS Translate lagged behind in translation quality, indicating reduced semantic fidelity and n-gram correspondence.
For language identification, no single model dominated across all metrics. AWS Translate achieved the highest accuracy (0.4717), making it more reliable for correct classifications overall. However, Claude attained the highest F1-score (0.7182) and recall (0.7581), suggesting its strength in capturing relevant instances. Gemini led in precision (0.7414), favoring correctness in its predictions. The best model depends on the specific performance criterion prioritized—accuracy, precision, or recall.
VI Discussion and Conclusion
Our evaluation of leading machine translation (MT) systems—ChatGPT, Claude, Gemini, Mistral, DeepL, AWS Translate, and LLaMA—on multilingual bug reports from the Visual Studio Code repository reveals both promise and limitations in domain-specific translation.
ChatGPT achieved the highest overall performance across all automatic metrics (BLEU, METEOR, ROUGE, BERTScore, and COMET), demonstrating its capability to retain both lexical fidelity and semantic integrity in technical contexts. Claude, Gemini, and Mistral also delivered competitive results, affirming that large language models (LLMs) can generalize well to structured, domain-specific inputs when properly tuned. In contrast, DeepL and AWS Translate underperformed on most translation quality metrics, possibly due to insufficient domain adaptation. Nevertheless, AWS Translate excelled in source language identification (accuracy: 0.4717, F1-score: 0.7027), while Claude and Mistral achieved higher F1-scores (0.7182 and 0.7142, respectively), suggesting trade-offs between precision and consistency.
Discrepancies among traditional metrics (BLEU, METEOR, ROUGE) and embedding-based metrics (BERTScore, COMET) emphasize the limitations of $n$-gram approaches in capturing semantic nuances. ChatGPT’s consistent top-tier performance across all five metrics challenges earlier assumptions about LLM shortcomings in technical translation and underscores the value of context-aware evaluation.
It is important to note the evolving nature of cloud-based MT systems. Regular updates to models such as ChatGPT and AWS Translate may lead to temporal variability in translation performance. Consequently, our results represent a time-bound snapshot rather than a static benchmark. Longitudinal studies are needed to track the progression and stability of these models over time.
While our dataset offers a realistic benchmark through bug reports from a popular open-source project, it remains limited to a single codebase. Expanding to include diverse projects and multilingual annotations would enhance generalizability. Additionally, reliance on a small group of domain-knowledgeable annotators introduces the potential for subjective bias.
In summary, modern MT systems—particularly LLM-based models—demonstrate strong potential for translating technical content like bug reports. Yet, challenges remain in preserving domain-specific terminology and structural fidelity. Future work should focus on improving neural architectures, enhancing domain adaptation techniques, and integrating specialized linguistic resources to support software engineering tasks.